OpenStudy (anonymous):

hey guys, i never really ask for help but im stuck in algebra 2. we had to make our own equations with information we gathered and i need help solving it. i reread the lesson a whole bunch of times but im still stuck. i need an explanation please!!!! <3

OpenStudy (anonymous):

my equation is a=26,000(1+ 2.38/12) ^(12)(5)

OpenStudy (anonymous):

i think its logarithms

OpenStudy (anonymous):

i really need help, if anyone has anything they could say to help would be great, ive been stuck on this the past 4 weeks >.<

OpenStudy (anonymous):

Algebra 2A or B?

OpenStudy (anonymous):

i dont know..i live in florida. we just have algebra 2 here..

OpenStudy (anonymous):

Ohh, idk im sorry :/

OpenStudy (anonymous):

:c do you have anyone you could refer to this? please im desperate. im now 4 weeks behind becaus eof one question

OpenStudy (anonymous):

I feel you, umm I could just message someone who had the name champion or mentor and they probably could help.. hope everything works out.

OpenStudy (anonymous):

that would be amazing..i fairly noob to this site

OpenStudy (phi):

can you take a screen shot of the original question, and post it?

OpenStudy (anonymous):

i made up the question...and i dont know how to use this site but i think i did it wrong it should be 26,000 = a(1+ 2.38/12) ^(12)(5)

OpenStudy (phi):

Can you fill in some details ? you said " we had to make our own equations with information we gathered" can you say what you info you used , and what the equation means ?

OpenStudy (anonymous):

it supposed to be like rate on a car loan. 2.38 is the rate and 26,000 is the price of the car, the 12 is per month and the 5 is over a time span of 5 years

OpenStudy (anonymous):

im supposed to find the amount paid after 5 years time

OpenStudy (anonymous):

using this formula A=P(1 + R/N) ^ (N)(T)

OpenStudy (phi):

ok. so you mean you borrow 26,000 at 2.38% interest rate for 5 years.

OpenStudy (anonymous):

um yes..i think

OpenStudy (anonymous):

26,000 is A 2.38% is R 12 months is N and 5 years is T

OpenStudy (phi):

the formula you posted A=P(1 + R/N) ^ (N)(T) is more appropriate to figure out how much money you have if you deposit 26,000 in the bank and it pays 2.38% interest for 5 years. A loan is more complicated. Does it have to be a loan ?

OpenStudy (anonymous):

thats the formula i HAVE to use and she just said that as an example i guess

OpenStudy (anonymous):

Thats what the formula should look like, just plug in the numbers

OpenStudy (phi):

ok. In that case, make the problem: You deposit 26,000 into a bank at 2.38% interest rate compounded monthly. How much money do you have after 5 years? there is one trick to using your formula. the rate r is a decimal. That means you must first change the 2.38% to 0.0238 before using it. P in your equation is how much you put in the bank at the start. P = 26000

OpenStudy (phi):

You should "tweak" your equation *** my equation is a=26,000(1+ 2.38/12) ^(12)(5) ***** to a=26,000(1+ 0.0238/12) ^(12)(5)

OpenStudy (anonymous):

ohhhhhhhh. i thought since its alreadya decimal i didnt have to change it????

OpenStudy (phi):

Usually interest rates are given as percents. Example: 10% which is the same as 0.10 or 5.5% which is 0.055 (divide by 100 to change the percent to a decimal)

OpenStudy (phi):

To answer the question How much money do you have after 5 years? use a calculator on your equation

OpenStudy (anonymous):

oh, ok. so i did the part in the parenthesis and got 1.00198333333

OpenStudy (anonymous):

do i take it to the 60th power?

OpenStudy (phi):

yes

OpenStudy (anonymous):

1.12623718056 now i have that

OpenStudy (phi):

yes. finally multiply that by the original principal 26000

OpenStudy (anonymous):

isnt this too big a number though??? 29282.1666947

OpenStudy (anonymous):

i thought it couldn't exceed 26,000?

OpenStudy (phi):

a=26,000(1+ 0.0238/12) ^(12)(5) a= 26000 (1.00198333333^60) a= 26000*1.12623718056 a= 29282.1666947 you should round that to the nearest penny

OpenStudy (anonymous):

so the extra is just what the bank makes then?

OpenStudy (phi):

if you put money in the bank they pay you interest (think of it as you are loaning the bank money and they pay you for it)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

how would i go about finding t?

OpenStudy (phi):

in other words, after 5 years, the 26,000 will grow to 29282.17 (rounded to the nearest cent)

OpenStudy (anonymous):

THANK YOU OH MY GOODNESS

OpenStudy (anonymous):

I struggled with this for weeks, muchos gracias

OpenStudy (phi):

to find "t" (I assume you mean the number of years) you need to know all the other numbers. Do you have a specific question where you want to find t ?

OpenStudy (anonymous):

i need to find how many months until id owe the bank $100,000

OpenStudy (anonymous):

so i would put that in for a and leave t blank?

OpenStudy (phi):

Can we change the question to how long before your bank account reaches 100,000 ? (loans normally have you start paying a monthly payment immediately, which is more complicated than your equation)

OpenStudy (phi):

can you write down what you think the equation should be, if you want to know how long it will take 26,000 to grow to 100,000 at an interest rate of 2.38% compounded monthly. ?

OpenStudy (anonymous):

yeah i guess

OpenStudy (anonymous):

100,000=26,000(1+ 0.0238/13)^(12)t

OpenStudy (phi):

I assume you mean divide by 12 (not 13) ?

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

woops

OpenStudy (phi):

the stuff inside the parens will be the same as before 1.00198333333

OpenStudy (anonymous):

ok then what?

OpenStudy (phi):

\[ 100,000=26,000(1.00198333333^{12t} )\] the next step is divide both sides by 26000

OpenStudy (anonymous):

3.84615384615

OpenStudy (phi):

you have an equation, so you should write the whole equation

OpenStudy (anonymous):

3.84615384615 = 1.001983333333^12t

OpenStudy (phi):

next, to get at the exponent, we use logarithms take the log of both sides (log base 10 is fine)

OpenStudy (phi):

we use this property on the right side \[ \log(a^b) = b \log(a) \]

OpenStudy (anonymous):

now im confused..

OpenStudy (phi):

you start with \[ 3.84615384615 = 1.001983333333^{12t}\] take the log of both sides \[ \log(3.84615384615 )= \log(1.001983333333^{12t})\]

OpenStudy (anonymous):

i got this for the left side 0.58502665202

OpenStudy (phi):

on the left side, you use a calculator. on the right side, you have to use the rule \[ \log(a^b) = b \log(a) \] for example \[ \log(1.01^{2t}) = 2t \log(1.01) \] on the left you get 0.58502665202 what about the right side ?

OpenStudy (anonymous):

and this for the right side 0.00086049767

OpenStudy (phi):

times 12 t

OpenStudy (anonymous):

0.01032597212t

OpenStudy (phi):

yes, but write out the whole equation 0.58502665202 = 0.01032597212t solve for t

OpenStudy (anonymous):

0.08333333268=t

OpenStudy (phi):

? try again 0.58502665202 = 0.01032597212t t= 0.58502665202/0.01032597212

OpenStudy (phi):

to solve for t we divide both sides by 0.01032597212

OpenStudy (anonymous):

sorry i have dyslexia >.< hold on

OpenStudy (anonymous):

56.6558426869=t

OpenStudy (phi):

I would round that to 56 and ⅔ years

OpenStudy (anonymous):

ok, thank you, SO MUCH. i fell so far behind in work and now i can finally move on.

OpenStudy (anonymous):

im online schooled and its hard for ME to teach myself, some peopel can do it but i have trouble, so again, thank you, LIFESAVER

OpenStudy (phi):

yw